# Bayesian Hierarchical Models for Quantitative Estimates for Performance metrics applied to Saddle Search Algorithms

### v1

@@ Abstract
The increasing use of high-throughput computational chemistry demands rigorous methods for evaluating algorithm performance. We present a Bayesian hierarchical modeling paradigm (brms/Stan) for analyzing key performance metrics: function evaluations, computation time, and success/failure. This framework accounts for variability across different systems and functionals, providing reliable uncertainty estimates beyond subjective visual assessments or frequentist limitations. We applied this to compare conjugate gradient (CG) and L-BFGS algorithms for the Dimer method’s rotation phase (EON, with/without removal of external rotations/translations) on a benchmark of 500 initial saddle search approximations, analyzing over 2000 runs. Our results show CG rotations generally outperform L-BFGS, exhibiting a statistically credible, small reduction in PES calls and significantly higher odds of successful convergence. Conversely, enabling rotation removal incurred a substantial PES call penalty without a corresponding credible improvement in success odds in this implementation. These findings, from our novel Bayesian hierarchical modeling application, suggest CG may be preferable for Dimer rotational optimization in similar contexts. This robust statistical framework highlights benefits for revisiting optimization strategies, quantifying uncertainty, and facilitating improved high-throughput computational chemistry methods.
@@ 4 Discussion
Counter-intuitively, enabling rotation removal, theoretically designed to simplify the optimization by projecting out external coordinates [27], resulted in a significant increase in computational cost (median +44% PES calls vs disabled) without a corresponding statistically credible improvement in success probability in our full model (OR for RotRem Yes vs No $\approx$ 1.9 [0.74, 5.07]). This discrepancy likely highlights the sensitivity of optimization performance to implementation details and the specific nature of the PES. The projection involved in rotation removal interacts with the optimizer’s steps and Hessian updates; for this dataset containing potentially fragmenting or aggregating systems near saddle points, where the separation of internal and external coordinates can be ill-defined [31], this projection may hinder convergence pathways that are accessible without it. Thus we observe less efficient trajectories when rotation removal is active for the majority of these systems.
This study also underscores the value of applying rigorous statistical modeling to computational benchmarks, moving beyond simple averages or qualitative comparisons. Traditional approaches can be easily confounded by the large system-to-system variability inherent in chemical datasets, as evidenced by the substantial standard deviation estimated for the random intercepts in our models i.e. $\sigma_{u}\approx$ 0.63 for PES calls (on the logarithmic scale), \$$\sigma$$\approx$\$3.6 for success odds(on the logit scale). Bayesian hierarchical models explicitly account for this variability through random effects ($u_{j}$). Furthermore, they leverage partial pooling (shrinkage), effectively borrowing strength across systems to produce more stable and reliable estimates for the effects of different methods, mitigating the influence of outlier systems - an effect sometimes colloquially related to “regression towards the mean”. This framework also provides direct quantification of uncertainty via full posterior distributions and credible intervals (e.g., Figs. 4, 5), facilitating more nuanced conclusions than point estimates or p-values alone [25, 26]. The ability to model interactions formally, although no strong interaction was detected here, is another advantage for dissecting performance factors.
There are very few results to imply why LBFGS for the rotation-phase would be of much use. The L-BFGS does not in most cases enjoy the same quadratic guarantee [3] as the BFGS, and additionally, the CG is less likely to be trapped by higher eigenpairs [24]. These results are not surprising in that respect, although it is clear that the difference is primarily in success probabilities, with no appreciable change in terms of computational performance.
@@ 5 Conclusions
We assert that these results, along with the automated capabilities of most workflow engines suggest a “chain” of methods approach rather than an absolute best. Although given the slight increase in performance and the significant increase in robustness it is hard to argue for the LBFGS rotations, the situation for the chemically intuitive external degree of freedom removal is more nuanced. We were able to comprehensively study the implementation of the rotation removal in EON. While the concept of removing external degrees of freedom is well-founded, our results indicate that, for the systems and implementation studied, this removal often increases the number of PES calls and reduces the likelihood of successful convergence, contrary to its intended purpose. There are however, situations where the rotation removal is required for convergence. Since the data demonstrates that for these random saddle point initializations that the rotation removal for such fragmenting systems is not optimal in most cases, it is more natural to first run CG-rotation without external rotation removal, and have a fallback rule to run CG-rotation with external rotation only for systems which otherwise fail. This strategy leverages the higher success rate of CG in the majority of cases while providing a fallback option for more challenging systems.
@@ 6 Supplementary Material
Model validations for the Bayesian models are fairly standard and verbose, so they are present in the accompanying supplementary material. The supplementary material also includes videos of the path taken to the saddle (or failure) for key systems referred to in the main text.

### v2

@@ Abstract
Rigorous performance evaluation is essential for developing robust algorithms for high-throughput computational chemistry. Traditional benchmarking, however, often struggles to account for system-specific variability, making it difficult to form actionable conclusions. We present a Bayesian hierarchical modeling framework that rigorously quantifies performance metrics and their uncertainty, enabling a nuanced comparison of algorithmic strategies. We apply this framework to analyze the Dimer method, comparing Conjugate Gradient (CG) and L-BFGS rotation optimizers, with and without the removal of external rotations, across a benchmark of 500 molecular systems. Our analysis confirms that CG offers higher overall robustness than L-BFGS in this context. While the theoretically-motivated removal of external rotations led to higher computational cost (¿40
@@ 4 Discussion
Counter-intuitively, enabling rotation removal, theoretically designed to simplify the optimization by removing global rotations of the coordinate vector [32], resulted in a significant increase in computational cost (median +44% PES calls vs disabled) without a corresponding statistically credible improvement in success probability in our full model (OR for RotRem Yes vs No $\approx$ 1.9 [0.74, 5.07]). This discrepancy likely highlights the sensitivity of optimization performance to implementation details and the specific nature of the PES. The projection involved in rotation removal interacts with the optimizer’s steps and Hessian updates; for this dataset containing potentially fragmenting or aggregating systems near saddle points, where the separation of internal and external coordinates can be ill-defined [36], this projection may hinder convergence pathways that are accessible without it. Thus we observe less efficient trajectories when rotation removal is active for the majority of these systems.
We underscore the value of applying rigorous statistical modeling to computational benchmarks, moving beyond simple averages or qualitative comparisons. Traditional approaches can be easily confounded by the large system-to-system variability inherent in chemical datasets, as evidenced by the substantial standard deviation estimated for the random intercepts in our models i.e. $\sigma_{u}\approx$ 0.63 for PES calls (on the logarithmic scale), \$$\sigma$$\approx$\$3.6 for success odds(on the logit scale). Bayesian hierarchical models explicitly account for this variability through random effects ($u_{j}$). Furthermore, they leverage partial pooling (shrinkage), effectively borrowing strength across systems to produce more stable and reliable estimates for the effects of different methods, mitigating the influence of outlier systems - an effect sometimes colloquially related to “regression towards the mean”. This framework also provides direct quantification of uncertainty via full posterior distributions and credible intervals (e.g., Figs. 4, 5), facilitating more nuanced conclusions than point estimates or p-values alone [30, 31]. The ability to model interactions formally is another advantage for dissecting performance factors.
Further theoretical considerations help to contextualize these empirical results. The rotational phase of the Dimer method represents a specialized optimization problem: isolating a single minimum-curvature mode. The Conjugate Gradient algorithm is inherently well-suited for such eigenvector-following tasks. It can also be less susceptible to being trapped by higher-energy eigenpairs during the search [29]. Conversely, while L-BFGS is a highly effective general-purpose optimizer, it does not, in most cases, share the same quadratic convergence guarantees as the full BFGS method [3]. This theoretical background aligns well with our findings, where the most significant performance gap between the two methods appeared in their relative robustness and success rates, rather than in the computational cost of already successful runs.
@@ 5 Conclusions
Our analysis of the Dimer method variants revealed that for this large benchmark set, the Conjugate Gradient (CG) optimizer for dimer rotation provided a clear advantage in convergence robustness over L-BFGS (Section 2.2). The decision to enable rotation removal is more complex. While theoretically motivated to simplify the potential energy surface, our findings indicate it led to higher computational costs for the majority of systems studied. However, our models also suggest that rotation removal may offer a reliability benefit when paired with the L-BFGS optimizer, highlighting a subtle interplay between algorithmic components that warrants further study.
We assert that these results, along with the automated capabilities of most workflow engines suggest a “chain” of methods approach rather than an absolute best. Although given the slight increase in performance and the significant increase in robustness it is hard to argue for the LBFGS rotations, the situation for the chemically intuitive external degree of freedom removal is more nuanced.
@@
There are however, situations where the rotation removal is required for convergence. Since the data demonstrates that for these random saddle point initializations that the rotation removal for such fragmenting systems is not optimal in most cases, it is more natural to first run CG-rotation without external rotation removal, and have a fallback rule to run CG-rotation with external rotation only for systems which otherwise fail. This strategy leverages the higher success rate of CG in the majority of cases while providing a fallback option for more challenging systems.
@@ 6 Supplementary Material
Model validations for the Bayesian models are fairly standard and verbose, so they are present in the accompanying supplementary material.
@@ Acknowledgments
Data Availability Statement
The full computational workflow, analysis code, and data supporting this study are openly available. A version-tagged GitHub repository (https://github.com/HaoZeke/brms_idrot_repro) provides all resources needed to reproduce the findings. To ensure full reproducibility, the repository contains a complete computational environment managed by pixi, with version-pinned software dependencies including EON, NWChem, and ASE, alongside utilities from Wailord [13]. In line with best practices for literate and reproducible programming [16, 17, 42, 15], some dependencies are bundled directly via gitsubrepo. For permanent access and citation, the final curated dataset is hosted on the Materials Cloud Archive [44] under the entry goswamiBayesianHierarchicalModels2025MCDat [14].